OpenStudy (goformit100):

Prove that the ten’s digit of any power of 3 is even. [e.g. the ten’s digit of 3^6 = 729 is 2].

OpenStudy (goformit100):

@TessaMarie08181997 @mikaela19900630

OpenStudy (rasperrystumper):

i dont get it explain more

OpenStudy (goformit100):

[e.g. the ten’s digit of 3^6 = 729 is 2].

OpenStudy (rasperrystumper):

ohh i get it okay let me see

OpenStudy (rasperrystumper):

none of them are

OpenStudy (anonymous):

did you try Induction proof?

OpenStudy (goformit100):

ya I tries but was unsuccessful using it

OpenStudy (anonymous):

\[\Large{3^1=03\quad 3^2=09\quad3^3=27\quad3^4=81\quad3^5=243\ldots\\ \text{we note that 1,3,9,7 are repeated cyclically every "4" steps of n }\\ 3^n=1\,{\rm or}\,3({\rm mod}\,10)\implies3^{n+1}=3k\,({\rm mod}\,10)\\ 3^n=9\,{\rm or}\,7({\rm mod}\,10)\implies3^{n+1}=3k\pm2\,({\rm mod}\,10) }\] we saw that k=2 for n=2 k=8 for n=3 by induction, this will be true for all positive powers of 3

OpenStudy (anonymous):

k = tens place digit

OpenStudy (goformit100):

Thankyou Sir